OpenStudy (anonymous):

Describe a pattern in the sequence of numbers. Predict the next number. 256, 16, 4, 2, . . .

jimthompson5910 (jim_thompson5910):

Notice how 16^2 = 256 4^2 = 16 and 2^2 = 4

OpenStudy (helder_edwin):

\[ \large a_1=256=16^2 \] \[ \large a_2=16=\sqrt{a_1} \]

OpenStudy (anonymous):

so whats the next number

jimthompson5910 (jim_thompson5910):

You take the square root of each term to get the next term Ex: sqrt(256) = 16